1.[A]the 18th term of an arithmetic progression is 25.find it 1st term if is common is 2? [b1]find the common ratio[b2]the 8th term [b3]nth term of the geometric progression
Answer by stanbon(75887)

1.[A]the 18th term of an arithmetic progression is 25.find it 1st term if is common is 2? [b1]find the common ratio[b2]the 8th term [b3]nth term of the geometric progression
======================
Comment: I'm not sure I understand your question.
#1. You have an arithmetic progression where a(18)=25 ; you want to find
the 1st term if the common difference is 2. If that is the question,
a(18) = a(1)+17d
25 = a(1) + 34
a(1) = -9
-------------------------
#2. Now you are talking about a geometric progression and you want to
find the ratio of the 8th term to the nth term. If that is the question
a(8)/a(n) = [1-r^9] /[1-r^(n+1)]
================
Cheers,
Stan H.